OpenStudy (anonymous):

What are the zeros and their multiplicities of the function f(x)=x^4-4x^3+3x^2?

OpenStudy (michele_laino):

please, start putting in evidence the common factor x^2, write yur answer

OpenStudy (anonymous):

How is the common factor x^2?

OpenStudy (cwrw238):

well x^2 will divide exactly into all 3 terms of the function

OpenStudy (anonymous):

\[ f(x)=x^4-4x^3+3x^2=x^2(x^2-4x+3)\]

OpenStudy (anonymous):

evidently \(0\) is a zero with multiplicity \(2\) as you have one factor of \(x^2\)

OpenStudy (anonymous):

the other zeros you can find by either factoring \(x^2-4x+3\) or using the quadratic formula